A camera lens used for taking close-up photographs 8
Answer #1

We know that the lens formula

1/f = 1/di + 1/do

1/22 = 1/di + 1/33

1/di = 1/22 -1/33 = 1/66

di = 66 mm............Ans. of a)

Magnification = di/do = 66/33 = 2.0 ........Ans. of (b)